OpenStudy (anonymous):

Find the magnitude & direction of the vector JK given points J(1,0) & K(-2,3).

OpenStudy (anonymous):

Magnitude = length of the line-segment = distance between J and K = sqrt( (1-(-2)) ^ 2 + (0 - 3)^2 ) (using distance between 2 points formula) = sqrt( 3^2 + 3^2 ) = sqrt(18) = 3 *sqrt(2) To get the direction, a quick way is: tan (angle) = (difference of y-coordinates) / (difference of x-coordinates) = (0 - 3) / (1 - (-2)) = -3/3 = -1 tan 45 = 1 but since we have -1, we subtract 45 from 180 to get 135 so direction is 135 from 'positive horizontal'. If you want, I could explain why we subtract from 180 if the tan is negative, and why use tan at all here.

OpenStudy (lgbasallote):

im interested to know why they subtract 180...i often get confused whether to subtract 180 or add 360 or whatever :/

OpenStudy (anonymous):

I think I get it now, thank you!(:

OpenStudy (anonymous):

welcome @shanti99